Question:
A college has seven instructors qualified to teach a special computer lab course which requires two instructors to be present. How many different pairs of teahers could there be?
(A) 42
(B) 49
(C) 15
(D) 21
Please help. Would the answer be (B) 49?
Thanks!

Answer by Cintchr(481)

This is like a factorial problem:
7! = 7 * 6 * 5 * 4 * 3 * 2 * 1 = 5040
BUT ... we are looking at only the first two spots. So
7 * 6 = 42